OpenStudy (anonymous):

lim of x approaching negative infinity (2x+3)/ (sqrt x^2+x+1) is...

OpenStudy (anonymous):

my guess is \(-2\)

OpenStudy (anonymous):

want me to check it?

OpenStudy (anonymous):

how would you get that answer?

OpenStudy (anonymous):

it could be wrong, because i did it with my eyeballs ignore everthing but the highest term \[\frac{2x}{\sqrt{x^2}}=\frac{x}{|x|}\] and since \(x<0\) you get \(-2\)

OpenStudy (anonymous):

typo there, should have been \[\frac{2x}{\sqrt{x^2}}=\frac{2x}{|x|}=-2\]

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

hold on maybe i am wrong lets check it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

you can also do some tedious calculation where you divide top and bottom by \(x\) but that is annoying and unnecessary

OpenStudy (anonymous):

ok thank you for your help!